### v1

# Interplay between topology and interactions in superconducting chains

Abstract
Most studies of non-trivial topological systems are carried out in non-interacting models that admit an exact solution. This raises the question, to which extent the consideration of electronic correlations, present in real systems, modify these results. Exact solutions of correlated electronic systems with non-trivial topological properties, although fundamental are scarce. Among the non-interacting soluble models, we single out the Kitaev p-wave superconducting chain. It plays a crucial role in clarifying the appearance of emergent quasi-particles, the Majorana modes, associated with non- trivial topological properties. Given the relevance of this model, it would be extremely useful if it can be extended to include correlations and still remain solvable. In this work we investigate a superconducting Kitaev chain that interacts through a Falicov-Kimball Hamiltonian with a background of localized electrons. For some relevant values of the parameters of this extended model solutions can be obtained, which allows for a detailed study of the interplay between electronic correlations and non- trivial topological behavior. Besides, the random occupation of the chain by the local moments brings the new interesting effects associated with disorder.
I Introduction
Systems with non-trivial topological properties have raised enormous interest in recent years [1, 2, 3, 4, 5]. They have expanded the landscape of interesting materials in the periodic table, giving rise to new concepts and to the discovery of unsuspected physical properties. Theoretical study of models with non-trivial topological properties has in many cases anticipated unusual behavior creating exciting challenges to experimentalists. One clear example is the Kitaev superconducting chain model, with emergent new quasi-particles, the Majorana fermions in its extremities [6]. The unusual properties of these modes, with possible applications in quantum computers [7, 8, 9, 10, 11, 12] is leading to enormous progress in experimental nano-physics [1, 3, 4, 5, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
Many of the models of topological superconductors and insulators are idealized, since they do not include electronic correlations, and consequently can be solved exactly. In one aspect this is very satisfactory, as it allows to obtain a clear picture of the nature of the topological non-trivial properties. On the other hand, it raises the question to which extent these properties are robust to electronic correlations present in real materials [24, 25, 26, 27]. The motivation of the present study is to explore this problem presenting a non-trivial topological interacting model with an exact solution for some values of its parameters. The model we focus in this work is the superconducting $p$-wave chain introduced by Kitaev, extended to include a coupling of the electrons to a background of localized electrons through a Falicov-Kimball (FK) interaction [28, 29]. This many-body problem can be mapped in a non-interacting system with a site dependent chemical potential, which has been extensively studied [31, 32, 32]. Some of these results can be directly used in our analysis, with peculiarities arising from the physical motivation of our model. Our work allows to explore on firm ground the interplay of electronic correlations and non-trivial topology. We consider different occupations of the sites of the chain by localized electrons, including the case of random occupations of these sites that give rise to new and interesting effects.
We organize the paper in the following way: In Sec. II, we introduce the Hamiltonian, which is a generalization of the standard Kitaev chain [6], including a background of localized electrons described by a Falikov-Kimball interaction [28]. In Sec. III, we present the formalism employed in calculating the exact solution of the model. In Sec. IV, we discuss the properties of the standard, spinless, one-dimensional FK model. We consider three cases: Homogeneous or uniform, staggered, and disordered configurations of the localized moments. In Sec. V, we consider the general case, where the conduction band has a p-wave superconducting pairing and interacts with the background of localized $f$-electrons. We present results of the density of states for the homogeneous and staggered cases. Finally, we study numerically the appearance of Majorana zero modes (MZMs) for the finite chain, including a simulation of the disordered case. Finally, in Sec. VI, we give a summary of the results, their physical consequences and new perspectives.
II The Hamiltonian
The Hamiltonian we consider, describes a chain of spinless, itinerant electrons with a $p$-wave pairing that interact with a background of localized electrons through a Falicov-Kimball interaction [28]. The Hamiltonian describing the $p$-wave superconductor is the Kitaev one [1, 6],
$\mathcal{H_{K}}=-\sum_{i,j}t_{ij}c^{\dagger}_{i}c_{j}-\sum_{i,j}\left(\Delta_{ij}c^{\dagger}_{i}c^{\dagger}_{j}+\Delta^{*}_{ij}c_{i}c_{j}\right)-\mu\sum_{i}c^{\dagger}_{i}c_{i},$ (1)
and that of the localized electrons [28],
$\mathcal{H_{L}}=\sum_{i}(E_{f}-\mu)f^{\dagger}_{i}f_{i}.$ (2)
In these equations, $t_{ij}$ is the hopping term of the conduction $c$-electrons, $\Delta_{ij}$ their $p$-wave coupling, $E_{f}$ the energy of the localized $f$-electrons and $\mu$ the chemical potential.
The coupling between the itinerant $c$-electrons and the localized $f$-electrons is of the FK type [28] and given by,
$\mathcal{H_{FK}}=J\sum_{i}\left(f^{\dagger}_{i}f_{i}-\frac{1}{2}\right)\left(c^{\dagger}_{i}c_{i}-\frac{1}{2}\right)=J\sum_{i}\left(n^{f}_{i}-\frac{1}{2}\right)\left(n^{c}_{i}-\frac{1}{2}\right).$ (3)
In order to solve the full Hamiltonian $\mathcal{H}=\mathcal{H_{K}}+\mathcal{H_{L}}+\mathcal{H_{FK}}$, we introduce sets of Majorana operators [35], $\alpha_{A}$, $\alpha_{B}$, $\beta_{A}$ and $\beta_{B}$, such that,
$c_{i}=(\alpha_{Bi}+i\alpha_{Ai})/2,$ (4)
$c^{\dagger}_{i}=(\alpha_{Bi}-i\alpha_{Ai})/2,$
where the Majoranas are their own anti-particles, i. e., $\alpha_{A,B}^{\dagger}=\alpha_{A,B}$. Similarly, for the $f$-electrons we have,
$f_{i}=(\beta_{Bi}+i\beta_{Ai})/2,$ (5)
$f^{\dagger}_{i}=(\beta_{Bi}-i\beta_{Ai})/2.$
In terms of these new operators we can write the interaction term as,
$\mathcal{H_{FK}}=\frac{J}{4}\sum_{i}(i\beta_{Bi}\beta_{Ai})(i\alpha_{Ai}\alpha_{Bi}).$ (6)
As for the other terms in the Hamiltonian, we get,
$\mathcal{H_{K}}+\mathcal{H_{L}}=-\frac{\mu}{2}\sum_{i}(1+\alpha_{Bi}\alpha_{Ai})-\frac{i}{2}\sum_{i}\left[(\Delta+t)\alpha_{Bi}\alpha_{Ai+1}+(\Delta-t)\alpha_{Ai}\alpha_{Bi+1}\right]-\frac{\mu-E_{f}}{2}\sum_{i}(1+\beta_{Bi}\beta_{Ai}).$ (7)
The many-body problem given by $\mathcal{H}=\mathcal{H_{K}}+\mathcal{H_{L}}+\mathcal{H_{FK}}$ is complex to solve, however a solution can be obtained in some cases.
III The solution
For further progress in the understanding the properties of the model, we consider the case where the chemical potential is at the energy of the localized particles, i.e., $\mu=E_{f}$ and for further simplicity, $E_{f}=\mu$ is taken to be zero. The case $E_{F}=\mu\neq 0$ also can be solved. Notice that for $\mu=0$, the superconducting chain, in the absence of the coupling to the localized electrons, is in a nontrivial topological state, since the condition $|\mu/2t|<1$ is satisfied [1]. On the other hand, for $\Delta=0$ the model corresponds to the one-dimensional spinless FK model [28].
When $E_{f}=\mu=0$, the Hamiltonian $\mathcal{H_{KL}}$ does not depend on the Majorana operators $\beta_{B}$ and $\beta_{A}$, associated with the localized electrons. We can then define an operator $K_{i}=i\beta_{Ai}\beta_{Bi}$ that commutes with the Hamiltonian of the superconductor, i.e., $[K_{i},\mathcal{H_{KL}}]=0$ for all sites. This implies that $K_{i}$ is a $c$-number [26]. On the other hand, the anti-commutation rules of the Majorana operators imply that $K_{i}^{2}=1$ or $K_{i}=\pm 1$. The Hamiltonian of the system can now be written as,
$\mathcal{H}=-\frac{i}{2}\sum_{i}\left[(\Delta+t)\alpha_{Bi}\alpha_{Ai+1}+(\Delta-t)\alpha_{Ai}\alpha_{Bi+1}\right]+i\sum_{i}J_{i}(i\alpha_{Ai}\alpha_{Bi}),$ (8)
where $J_{i}=JK_{i}/16$ ($K_{i}=\pm 1$). From a formal point of view $\mathcal{H}$ is a quadratic form that can be exactly solved. The interacting problem has been mapped in a non-interacting one, formally in a system with a site dependent chemical potential [31, 32, 33].
Let us introduce the quantities
$\alpha_{Ai}=\sum_{k}e^{ikr_{i}}\alpha_{A}(k),$ (9)
and an equivalent equation for $\alpha_{Bi}$. Since $\alpha_{A}^{\dagger}(k)=\alpha_{A}(-k)$, only $\alpha_{A}(k=0)$ is a Majorana operator [35]. In terms of these operators the Hamiltonian can be written as,
$\mathcal{H}=-\frac{i}{2}\sum_{k}\left[(\Delta+t)e^{ik}\alpha_{B}(k)\alpha_{A}(-k)-(\Delta-t)e^{-ik}\alpha_{B}(k)\alpha_{A}(-k)\right]+iJ\sum_{k}\alpha_{A}(k)\alpha_{B}(-(k+Q)),$ (10)
where we wrote $J_{i}=Je^{iQr_{i}}$. The homogeneous, or uniform configuration, with a localized electron on each site of the lattice, $n_{f}=1$, corresponds to $Q=0$. The staggered configuration, where the values of $K_{i}$ alternate between ±$\pm 1$, such that $n_{f}=1$ and $n_{f}=0$, in consecutive sites, corresponds to $Q=\pi$ ($n_{f}=(1/2)(1+i\beta_{Ai}\beta_{Bi})$). Among all possible configurations, these two play an important role, as we discuss below.
The excitations of the system can be obtained from the Hamiltonian of the infinite chain, Eq. 10. For the uniform configuration ($Q=0$), we get,
$\omega_{h(1,2)}(k)=\frac{\sqrt{\Delta^{2}+2J^{2}\pm 4Jt\cos(k)+\cos(2k)\left(t^{2}-\Delta^{2}\right)+t^{2}}}{\sqrt{2}},$ (11)
and $\omega_{h(3,4)}(k)=\pm\omega_{h(1,2)}(k)$. The excitations in the staggered configuration with $Q=\pi$ are given by
$\omega_{s(1,2)}(k)=\frac{\sqrt{\Delta^{2}+2J^{2}\pm 4\Delta J\sin(k)-\cos(2k)\left(t^{2}-\Delta^{2}\right)+t^{2}}}{\sqrt{2}},$ (12)
and $\omega_{s(3,4)}(k)=\pm\omega_{s(2,1)}(k)$. The condition $\mu=E_{F}=0$ implies that for the non-interacting system, the band of conduction electrons corresponds to a half-filled band of (spinless) conduction electrons. The number of localized $f$-electrons is arbitrary, $n_{f}\leq 1$, and does not affect the total energy of the system in the non-interacting case, $J=0$.
In Fig. 1 we plot the dispersion relations, for specific values of parameters, of the uniform and staggered configurations. For completeness we also show the dispersion of the pure Kitaev chain ($J=0$).
Figure 1: (Color online) Dispersion relations of the excitations in the superconducting state for the a) homogeneous case, $\omega_{h}(k)$, and b) staggered case, $\omega_{s}(k)$, for $t=1$, $\Delta/t=0.2$ and $J/t=0.9$. c) energy of the bogoliubons $\omega_{0}(k)$, of the non-interacting case with the same parameters, but $J=0$ (Kitaev chain).
In Section IV, we will discuss the nature of the superconducting phases in the presence of the interaction with the localized electrons. We investigate how this interaction modifies the topological properties of the Kitaev chain. We performed simulations in finite chains to study the robustness of the Majorana modes in the Kitaev chain to the coupling to the localized fermions. These simulations also yield the ground state energy of different configurations and allow to identify the ground states.
However, before going deeper in the study of the superconducting states, we discuss briefly the spinless FK model in electronic systems. In our model this corresponds to taking the superconducting gap $\Delta=0$ in Eq. 8.
IV The one-dimensional spinless Falicov-Kimball model ($\Delta=0$)
The FK model is well studied in strongly correlated electronic systems, with many applications ranging from metal-insulator to valence transitions [28]. In particular the one-dimensional, spinless FK model has many well-known properties [28]. It can be approached using the technique of bosonisation [28] that yields exact results [28]. From the point of view of topological properties this model is completely trivial.
In this section we briefly discuss the spinless one-dimensional FK model. Formally, we just take the amplitude of the superconducting gap $\Delta=0$ to obtain the dispersion relations for this model. We can distinguish two situations depending on the value of the ratio $(J/t)$.
Figure 2: a) b) c), dispersion relations $\omega_{h}(k)$ of the excitations of the FK model ($t=1$ and $\Delta/t=0$), for the homogeneous case for $J/t=0.75$, $1$ and $1.2$, respectively. d) e) f), dispersion relations $\omega_{s}(k)$ for the staggered case for $J/t=0.75$, $1$ and $1.2$, respectively. In the homogeneous case a) corresponds to a metallic state and at $J/t=1$, (Fig. 2b) there is a metal insulator transition. In the staggered configuration the system is always insulating, unless $J=\Delta$, where it is a Dirac semi-metal. The black line is the energy of the localized quasiparticles that coincides with the chemical potential.
Among the special configurations, we distinguish two cases. For $J/t>1$, both homogeneous and staggered configurations correspond to insulating phases as shown in Fig. 2. Conversely, for $J/t<1$, the homogeneous configuration corresponds to a metallic phase, but the staggered one remains insulating, as can be seen in Fig. 2. The homogeneous configuration corresponds to one localized $f$-electron per site and the staggered to $1/2$ localized electron per site. Since we are working in the grand canonical ensemble, with the condition $E_{f}=\mu=0$, the total energy of the system does not depend directly on the number of $f$-electrons. The dependence on $n_{f}$ is only indirect, since the total energy, given by that of the $c$-electrons is determined by their coupling to the former. In Fig. 3 we show the ground state energies of the special configurations, $Q=0$ and $Q=\pi$, as functions of the interaction $J$. Notice that the staggered configuration has a lower energy compared to the uniform one. In the homogeneous case, with one localized electron per site ($n_{f}=1$), there is a metal-insulator transition at $|J/t|=1$, as shown in Fig. 2 for the dispersion relations.
We point out that the staggered configuration is not the only possibility for a number of localized electrons per site, $n_{f}=1/2$. Unlike the $Q=0$ uniform configuration, there are many other configurations with $<n_{f}>=1/2$, besides the $Q=\pi$, staggered one [29]. These configurations, which are mainly disordered, may become important at high temperatures, as they increase the entropy of the system [28, 29]. In Fig. 3 show the ground state energies, as functions of the interaction $J$, of the special configurations and of random configurations with $<n_{f}>=1/2$. In these disordered configurations there is a probability $p=1/2$ for a site being occupied, or not, by a localized $f$-electron. Notice that the ordered, $Q=\pi$, staggered configuration is always the one with lower energy. Our choice of $\mu=E_{f}=0$ implies that for spinless electrons there is and average number $<N_{c}>=1/2$ of $c$-electrons per site and an arbitrary number of localized $f$-electrons ($n_{f}\leq 1$). The uniform and staggered configurations fixes the number of localized electrons at $n_{f}=1$ and $n_{f}=1/2$, per site, respectively. The former case has a metallic ground state and the latter is an insulator. This closely resembles the Peierls instability where there is a doubling of the unit cell as the system goes from a homogeneous to a staggered configuration [30].
Figure 3: (Color online) Ground state energy of the spinless FK model ($\Delta=0$). Homogeneous case, $n_{f}=1$ (red), staggered, $n_{f}=1/2$ (blue) and random configurations (green) with $<n_{f}>=1/2$. Notice that, the staggered configuration has an energy lower then all the probed random configurations with $<n_{f}>=1/2$.
V Topological superconductivity in the presence of Interaction between itinerant and localized electrons
We now consider the case where the conduction band has a $p$-wave superconducting pairing and interacts with the background of localized $f$-electrons. The simple, non-interacting case with $J=0$, corresponds to a Kitaev chain and an independent collection of localized electrons. The topological properties of the Kitaev chain are well known, it presents a non-trivial topological phase, for $|\mu/2t|<1$, with Majorana modes at its ends [1]. In the present case the chemical potential $\mu=0$, such that the Kitaev chain is in a topological superconducting phase [1]. We now turn on the interaction with the localized $f$-electrons and study its effect on the topological properties of the chain using simulations on finite chains.
Figure 4: Plot of the density of states in the superconducting state. a) for the uniform case for different values of the coupling parameter $J$, below, at, and above the topological transition. b) for the uniform case at the topological transition, for different pairing amplitudes $\Delta$. c) for the staggered configuration as a function of $J$ for fixed $\Delta$.
Figure 5: (Color online) Eigenvalues $\lambda/t$ for a chain of 300 sites for $t=1$, $\Delta/t=0.2$, a) in the uniform phase, for $J/t=0.1<1$, b) in the staggered phase for $J/t=0.1<\Delta/t$. In c) and d) we show the amplitude square of the wave functions of the zero modes, i.e, eigenvalues closest to zero, for the uniform and staggered phases, ( a) and b) above), respectively.
Figure 6: (Color online) Eigenvalues $\lambda/t$ for a chain of 300 sites for $t=1$, $\Delta/t=0.2$, a) in the uniform phase, for $J/t=1.3>1$, b) in the staggered phase for $J/t=1.3>\Delta/t$. Both phases are gapped and topological trivial superconductors as indicated by the absence of zero energy surface modes.
Figure 7: (Color online) Ground state energy of the $p$-wave superconductor as a function of the interaction $J$ for $\Delta/t=0.5$. Uniform case, $n_{f}=1$ (red), staggered, $n_{f}=1/2$ (blue) and 200 random configurations (green) with $<n_{f}>=1/2$. The staggered configuration has an energy lower than all the probed random configurations with $<n_{f}>=1/2$.
Figure 8: a) Eigenvalues $\lambda/t$ for $\Delta=0.2$, corresponding to 200 realizations of random configurations, with probability $p=1/2$ for a site being occupied or unoccupied by a localized electron ($<n_{f}>=1/2$) for a chain with 40 sites. a) for $J=0.1$, b) for $J=0.3$, c) for $J=1$ and d) for $J=1.2$ The zero modes persist all the way to the topological phase transition of the uniform case at $J/t=1$.
First, we discuss the results of Fig. 1 for the dispersion relations of the infinite Kitaev chain coupled to the local moments. We consider both, the uniform and staggered configurations, for some relevant set of parameters. We take, $t=1$, so that the hopping sets the energy scale for the problem, and for the pairing gap we take $\Delta/t=0.2$. In the uniform case, for $J/t=1$ there is a gap closing, and the system is a gapless superconductor. For $J\neq 1$ the superconductor is gapped, as can be seen directly from the dispersion relations and density of states in Figs. 4a. However, there is a fundamental difference between the cases $J/t<1$ and $J/t>1$, as we discuss below for finite chains.
Figure 9: a) Amplitude square of the wave functions of the lower energy modes of a chain with 180 sites for a particular realization of disorder. a) $J=0.1$. b) $J=0.2$. c) Amplitude of the wave-function of the ordered staggered configuration with $J=0.1$. Compare with a) and notice that in the random case the wave function is more localized.
Fig. 4b shows the density of states for the homogeneous case in the gapless case ($J=1$), as a function of the $p$-wave superconductor parameter $\Delta$. For $\Delta$ close to zero, the FK strong correlations produce a huge peak at $\mu=0$. As the pairing interaction is turned on, a competition between the localization due to FK interactions and the formation of $p$-wave pairs occurs; the peak decreases as the superconducting pairs are formed, as indicated in the inset of the figure, and the density of states approaches that of a renormalized uncorrelated tight-binding chain. For comparison, we also show a tight-binding density of states for $J=\Delta=0.0001t$.
In the staggered configuration, gapless superconductivity occurs at $J=\Delta$ and again the superconductor is gapped whenever $J\neq\Delta$ as shown in Fig. 4c. Notice in Fig. 4c that there is a finite density of states at around $\mu=0$ for $J=\Delta$, i.e., at the topological transition of the staggered configuration where the system is gapless. This is similar to the corresponding density of states of an armchair nano-ribbon, whenever $N=3M-1$, with $M$ being integer [34]. As in the uniform case there is an essential difference between the gapped phases, in the staggered phase for $J>\Delta$ and $J<\Delta$, as concerns their topological properties.
In order to investigate the topological properties of the interacting system, instead of calculating a topological invariant [31, 32, 33] and using the bulk-boundary correspondence principle [34], we directly investigate the behavior of interacting finite chains using open boundary conditions. In this approach the appearance of zero energy surface states gives direct evidence of the non-trivial topological character of the system.
In Fig. 5 we show the eigenvalues for a chain of 150 sites for both uniform and staggered configurations. In the case of the uniform chain, the superconductor is gapped with zero energy modes that persist, as long as $J<1$. These zero modes are surface states as can be seen directly from their wave functions shown in the same figure. At $J=1$ the gap closes and the system is a gapless superconductor. For $J>1$, the gap reopens, as shown in Fig. 1, but the zero energy surface states disappear, indicating that the superconductor is topologically trivial. Notice that in the absence of pairing, the FK chain is metallic for $J<1$, as shown in Fig. 2a. These results show that in the uniform case there is a topological phase transition, with a gap closing, at $J/t=1$.
In the staggered case, the topological superconducting phase, characterized by the presence of zero energy surface states, exists for $J<\Delta$, as shown in Fig. 5b. At $J=\Delta$ the superconductor is gapless and above this critical value, the superconducting chain is gapped and topologically trivial with no zero modes. In this staggered case, as we turn on the superconducting pairing, non-trivial topological superconductivity emerges from an insulating phase (see Fig. 2d of the FK model, as soon as, the pairing amplitude becomes larger than $J$. Then, the topological phase transition occurs for $J/t=1$ in the uniform case and for $J/\Delta=1$ in the staggered case.
In Fig. 6, we show the eigenvalues of a finite chain for values of the coupling $J$ above the topological transition, both in the uniform and staggered configurations. Notice the absence of zero modes in the gap showing that these phases are topologically trivial.
For the uniform case, with a localized electron at each side of the chain there is no room for disorder. This is different in the staggered case, with $n-f=1/2$, where the localized electrons may also be randomly distributed in the chain. In Fig. 7 we plot the ground state energies of the superconductor for the uniform, staggered and random configurations as a function of the coupling $J$. Notice that for $n_{f}=1/2$ the ordered staggered configuration has always the lower energy and is the true ground state for $<n_{f}>=1/2$.
In the random case, with $<n_{f}>=1/2$, a given site has a probability $p=1/2$ of being empty or occupied by a local moment. It is interesting to explore the topological character of the random chains [37, 38, 23]. Fig. 8 shows the eigenvalues for a chain with 40 sites and 200 realizations of disorder. In this case we assume that a given site has a probability $p=1/2$ of being empty or occupied by a local moment. This corresponds to taking $<n_{f}>=1/2$. We observe the presence of two zero energy surface states on average for the whole range of couplings, $J\leq 1$, as shown in Fig. 8. The wave functions of these zero modes are shown in Fig. 9 showing that these are really surface states. Compared to the ordered staggered configuration, where topological superconductivity occurs for $J_{c}=\Delta$, we conclude that disorder extends the range of topological superconductivity to much larger values of the coupling. This is a surprising effect that has been observed previously in similar systems [23].
The critical value $(J/t)_{c}=1$ that we observe numerically for the disordered system with $p=1/2$, is the same as for the homogeneous case and different from the ordered staggered configuration with the same average number of localized electrons per site, $<n_{f}>=1/2$. Notice that can write the interaction between localized and conduction electrons at a given site $i$ of the chain, as $J_{i}=Jx_{i}$ where $x_{i}=\pm 1$. Consider a random occupation of the local moments in the chain, given by, $P(x_{i})=p\delta(x_{i}-1)+(1-p)\delta(x_{i}+1)$ for a given site $i$. Since the $x_{i}$s are independent random variables the critical value $(J/t)_{c}=1$ for the topological transition is a consequence of the invariance of the Hamiltonian under the exchange $J\rightarrow-J$.
VI Conclusions and perspectives
In this work, we have studied the properties of a $p$-wave superconducting chain in the presence of an interaction between the conduction electrons and a background of localized electrons. The interaction is of the FK type and couples the charge degrees of freedom of both systems, itinerant and localized electrons. We have shown that the many-body problem can be mapped in a non-interacting one with a site dependent chemical potential, which allows it to be soluble. The solution provides a clear picture of the effects of interactions in the topological properties of the chain. We find that in the case of a chain coupled to a uniform configuration of localized electrons, i.e., with every site of the chain occupied by a localized electron, the interaction between the two systems leads eventually to the suppression of the topological phase. However, the critical value of this coupling is large, of the order of the bandwidth of the conduction electrons. In the case the localized electrons are in a staggered ordered configuration, in which every other site in the chain is occupied, the critical value of the coupling to suppress the topological phase is much reduced, being given by the pairing energy of the conduction electrons. In this case superconductivity arises from an insulating state of the FK aboce a critical value of the pairing amplitude, $\Delta_{c}=J$. This staggered configuration corresponds to $<n_{f}>=1/2$, and for most of the configurations with this concentration, the localized electrons randomly occupy the sites in the chain. We have numerically studied the case of random configurations with probability $p=1/2$ for a site in the chain being occupied or not by a localized electron, which corresponds to $<n_{f}>=1/2$. We find that randomness extends the range of stability of the topological superconducting phase with respect to the coupling to the localized electrons. While in the ordered staggered configuration, the topological superconducting phase exists for $J<\Delta$, in the presence of disorder it is stable up to $J=t$, the critical coupling of the uniform configuration. We find however that the ordered staggered configuration has always the lower energy. The exact results of the model allows to put on firm grounds the role of interaction and disorder in the topological properties of an archetypal system, namely the Kitaev chain. The model, can be extended to take into account different occupations for localized electrons with $p\neq 1/2$. Also, we can assume that the energy of the localized electrons coincide with the chemical potential, i.e., $E_{f}=\mu$, but relax the condition in the chemical potential so that $\mu\neq 0$. In this case we can study the phase diagram for different occupations of the band of conduction electrons.
An interesting challenge is to include hybridization between the localized and itinerant electrons, which gives rise to the quantum Falicov-Kimball model [28].
We could also investigate the thermoelectric properties of those systems at the topological quantum phase transition of the homogeneous ($t=\Delta$) and staggered ($J=\Delta$, cases [39]. Another possibility is to extend include spin in the model and study the spin polarization of MZMs in nanoribbons [18].
Acknowledgements.
M A C acknowledges Brazilian National Council for Scientific and Technological Development (CNPq) Grant Number 305810/2020-0 and Foundation for Support of Research in the State of Rio de Janeiro (FAPERJ), Grant Number 201223/2021 for partial financial support. M S F acknowledges financial support from the National Council for Scientific and Technological Development (CNPq) Grant Number 311980/2021-0 and from the Foundation for Support of Research in the State of Rio de Janeiro (FAPERJ) process number 210 355/2018. A C P L acknowledges, Coordenação de Aperfeiçoamento do Ensino Superior (CAPES) for financial support.